A group of 72 children and adults attended a soccer game on Tuesday. Each adult paid $5 and each child paid $2. The group paid a total of $264. Write and solve a system of equations for this problem.

Answer by addingup(3677)

Solving by substitution:
a+c = 72 subtract a from both sides:
a = 72-c
-------------------------------------------
5a+2c = 264 substitute for a
5(72-c)+2c = 264
360-5c+2c = 264
-3c = -96
c = 32
-----------------------------------------
Solve by elimination:
a+c = 72 multiply all times -5 and add to the next equation
-------------------------------
5a+2c = 264
+
-5a-5c = 360
0a-3c = -96
-3c = -96
c = 32